### v1

@@ 4.1.4 Meta-generalization to deeper networks
Figure 6: Evaluating generalization capabilities of $\mu\text{LO}_{S}$ to deeper networks. The figures report the performance of learned optimizers for training depth-16 ViTs for image classification, Transformers for language modeling, and MLPs for image classification. We plot average training loss over $5$ seeds with standard error bars.
@@
Figure 7: Evaluating generalization capabilities of $\mu\text{LO}_{S}$ to longer training horizons. $\mu\text{LO}_{S}$ is meta-trained for $1000$ steps of training (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We observe that the optimizer seamlessly generalizes to training horizons much longer than meta-training (25$\times$ longer, here).
@@ 4.1.5 Meta-generalization to longer training horizons
In this subsection, we evaluate the capability of $\mu$LOs to generalize to much longer training horizons than those seen during meta-training. Specifically, we use $\mu\text{LO}_{M}$ and $\text{LO}_{M}$ to train a $w=1024$ 3-layer MLP and ViT on $32\times 32\times 3$ ImageNet and a $w=1024$ 3-layer Transformer for autoregressive language modeling on LM1B. Each model is trained for $25,000$ steps (25$\times$ the longest unroll seen at meta-training time). Figure 7 reports the training loss averaged over $5$ random seeds. We observe that $\mu\text{LO}_{M}$ stably decreases training loss over time for each task, while $\text{LO}_{M}$ fails to decrease training loss (a), decreases it but suffers instabilities (b), or diverges after $8000$ steps (c). These results suggest that generalization to longer training horizons is another benefit of using $\mu$P with learned optimizers.
@@ 5 Limitations
While we have conducted a systematic empirical study and shown strong results within the scope of our study, there are a number of limitations of our work. Specifically, (1) we do not meta-train on tasks other than MLPs for image classification and we do not provide evaluation of models wider than $8192$ (MLPs) and $2048$ (transformers) due to computational constraints in our academic environment.
@@ 6 Conclusion
We have demonstrated that parameterizing networks in $\mu$P significantly improves the generalization of LOs to wider networks than those seen during training, even surpassing VeLO (meta-trained for $4000$ TPU months) on the widest tasks. Moreover, our experiments show that networks trained with $\mu$P generalize better to wider and deeper out-of-distribution tasks than their SP counterparts. When evaluated on much longer training tasks, we observe that $\mu$P has a stabilizing effect, enabling meta-generalization to much longer unrolls (25×25$\times$25 × maximum meta-training unroll length). All of the aforementioned benefits of $\mu$LOs come at zero extra computational cost compared to SP LOs. Our results outline a promising path forward for low-cost meta-training of learned optimizers that can generalize to large unseen tasks.
@@ Acknowledgements
Appendix A Table of Contents
Appendix B Proof of proposition 1
@@
The update produced by $f_{\phi}$ is denoted $\Delta W$ and we write $\nabla W$ the corresponding gradient, so that $\Delta W=f_{\phi}(\nabla W)$. For the sake of simplicity, $n$ will be the output size and $d$ the feature input size of our neural network. Our goal is to satisfy the desiderata of [43, Appendix J.2]. We assume our initialization follows Initialization-$\mu$ in Sec 3. Overall, our goal is to study strategy so that if $x_{i}=\Theta(1)$, then one needs to renormalize/initialize so that $(Wx)_{i}=\Theta(1)$ while $((W+\Delta W)x)_{i}=\Theta(1)$ so that the update is as large as possible. Note that given the assumptions on $f$, if $x=\Theta(\frac{1}{n})$, then $f(x)=\Theta(1)$.
@@
Appendix C Hyperparameter tuning $\mu$Adam
We tune two $\mu$Adam baselines on the same meta-training tasks as our learned optimizers, creating analogous baselines. $\mu$AdamS is tuned on an ImageNet classification task with a 3-layer $128$ width MLP, while $\mu$AdamM was tuned using a $1024$ width MLP for the same task. As mentioned in section 4, for each task, we tune the learning rate and three multipliers: input multiplier, output multiplier, and the hidden learning rate multiplier. These multipliers correspond to adding a tunable constant to the pre-activation multiplier for input weights, the pre-activation multiplier for output weights, and the Adam LR for hidden weights (e.g., in Table 8 of [43]). Specifically, we search for the learning rate in ${0.1,0.01,0.001,0.0001}$ and for each multiplier in $\{2^{-4},2^{-2},1,2^{2},2^{4}\}$. This results in a grid search of More details about the grid search over $500$ configurations, whose optimal values are reported in table 2.
Table 2: Best hyperparameters values for $\mu$Adam baselines. Both optimizers are tuned to optimize and MLP on $32\times 32\times 3$ ImageNet classification. However, $\mu$AdamS optimizers a 3-layer W$=128$ MLP, while $\mu$AdamS optimizers a 3-layer W$=1024$ MLP.
@@ Appendix C Hyperparameter tuning $\mu$Adam
Appendix D Meta-training with $\mu$LOs
General meta-training setup
@@ Appendix D Meta-training with $\mu$LOs
Each learned optimizer is meta-trained for $5000$ steps of gradient descent using AdamW [20] and a linear warmup and cosine annealing schedule. We using PES [36] to estimate meta-gradients with a truncation length of $50$ steps and sampling $8$ perturbations at each step with standard deviation $0.01$. For the inner optimization task, we used a maximum unroll length of $1000$ iterations; that is, all our learned optimizers see at most $1000$ steps of the inner optimization problem during meta-training. Unlike with $\mu$Adam, we do not tune the $\mu$P multipliers when meta-training $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$, instead, we set the all to $1$. All optimizers are meta-trained on a single A6000 GPU. $\mu\text{LO}_{S}$ and $\text{LO}_{S}$ take $8$ hours each to meta-train, while $\mu\text{LO}_{M}$ and $\text{LO}_{M}$ take $103$ hours.
Ablating meta-training hyperparameters for $\mu$LOs
@@
While there are very few differences between $\mu$LOs and SP LOs, the effective step size for hidden layers is changed (see eq. 3) which could alter the optimal meta-training hyperparameters. Consequently, we conduct an ablation study on hyper-parameters choices for $\mu\text{LO}_{S}$. Specifically, using AdamW and gradient clipping with a linear warmup and cosine annealing LR schedule, we meta-train $\mu\text{LO}_{S}$ to train 3-layer width 128 MLPs to classify $32\times 32\times 3$ ImageNet Images. By default, we warmup linearly for 100 steps to a maximum learning rate of $33e-$ 3 and anneal the learning rate for $4900$ steps to a value of $31e-$ 3 with $\lambda_{1}=0.001$ (from equation 3) and sampling $8$ perturbations per step in PES[36]. The above ablation varies the maximum learning rate $\in\{1e-2,3e-3,1e-3\}$ (always using $100$ steps of warmup and decaying to 0.3×0.3$\times$0.3 ×MaxLR), $\lambda_{1}\in\{0.001,0.01,0.1\}$, the number of steps (5k or 10k), and the number of perturbations (8 or 16). We observe that using all default values except for $\lambda_{1}=0.01$ yields one of the best solutions while being fast to train and stable during meta-training. We, therefore, select these hyperparameters to meta-train $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$.
@@
While we use the same $\mu$P at meta-training and testing time, it is important to consider meta-training tasks that have similar training trajectories to their infinite width counterparts. In [43], authors provide discussions of these points for zero-shot hyperparameter transfer. Two notable guidelines are to initialize the output weight matrix to zero (as it will approach zero in the limit) and to use a relatively large key size when meta-training transformers. For all our tasks, we initialize the network’s final layer to zeros. While we do not meta-train on transformers, we suspect that the aforementioned transformer-specific guidelines may be useful.
Figure 8: Ablating Meta-training Hyperparameter for $\mu\text{LO}_{S}$. Using AdamW with a linear warmup and cosine annealing schedule, we meta-train $\mu\text{LO}_{S}$ to train 3-layer width 128 MLPs for classifying $32\times 32\times 3$ ImageNet Images. By default, we warmup linearly for 100 steps to a maximum learning rate of $33e-$ 3 and anneal the learning rate for $4900$ steps to a value of $31e-$ 3 with $\lambda_{1}=0.001$ (from equation 3) and sampling $8$ perturbations per step in PES[36]. The above ablation varies the maximum learning rate $\in\{1e-2,3e-3,1e-3\}$ (always using $100$ steps of warmup and decaying to 0.3×0.3$\times$0.3 ×MaxLR), $\lambda_{1}\in\{0.001,0.01,0.1\}$, the number of steps (5k or 10k), and the number of perturbations (8 or 16). We observe that using all default values except for $\lambda_{1}=0.01$ yields one of the best solutions while being fast to train and stable during meta-training.
Appendix E Features of the learned optimizer
Table 3: Per-parameter features used as input to our learned optimizers. All the coefficients, $\beta_{i}$, are learnable parameters adjusted during meta-optimization. We replicate the table of [15] for convenience.
@@ 3 $m$ Adafactor normalized values $m_{t,i}\times\textsc{row factor}\times\textsc{column factor}$
Appendix F Extended Results
F.1 List of Meta-testing Tasks
Table 4 reports the configuration of different testing tasks used to evaluate our optimizers. We note that we do not augment the ImageNet datasets we use in any way except for normalizing the images. We tokenize LM1B using a sentence piece tokenizer[18] with 32k vocabulary size. All evaluation tasks are run on A6000 or A100 GPUs for $5$ random seeds. Each individual run takes less than $2$ hours to complete on a single 80GB A100.
@@ F.1 List of Meta-testing Tasks
Identifier Dataset Model Depth Width Batch Size Sequence Length ${\mathcal{T}}^{\text{MLP}}_{1}$ $32\times 32$ ImageNet MLP $3$ $128$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{2}$ $32\times 32$ ImageNet MLP $3$ $512$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{3}$ $32\times 32$ ImageNet MLP $3$ $2048$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{4}$ $32\times 32$ ImageNet MLP $3$ $8192$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{5}$ $64\times 64$ ImageNet MLP $3$ $128$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{6}$ $64\times 64$ ImageNet MLP $3$ $512$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{7}$ $64\times 64$ ImageNet MLP $3$ $2048$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{8}$ $32\times 32$ Cifar-10 MLP $3$ $128$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{9}$ $32\times 32$ Cifar-10 MLP $3$ $512$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{10}$ $32\times 32$ Cifar-10 MLP $3$ $2048$ $4096$ ${\mathcal{T}}^{\text{MLP}}_{11}$ $32\times 32$ Cifar-10 MLP $3$ $8192$ $4096$ ${\mathcal{T}}^{\text{LM}}_{1}$ LM1B Transformer LM $3$ $192$ $128$ $64$ ${\mathcal{T}}^{\text{LM}}_{2}$ LM1B Transformer LM $3$ $1024$ $128$ $64$ ${\mathcal{T}}^{\text{LM}}_{2}$ LM1B Transformer LM $3$ $2048$ $128$ $64$ ${\mathcal{T}}^{\text{ViT}}_{1}$ $32\times 32$ ImageNet ViT $3$ $192$ $1024$ ${\mathcal{T}}^{\text{ViT}}_{2}$ $32\times 32$ ImageNet ViT $3$ $1024$ $1024$ ${\mathcal{T}}^{\text{ViT}}_{2}$ $32\times 32$ ImageNet ViT $3$ $2048$ $1024$ ${\mathcal{D}}{\mathcal{T}}^{\text{MLP}}_{1}$ $32\times 32$ ImageNet MLP $16$ $512$ $4096$ ${\mathcal{D}}{\mathcal{T}}^{\text{ViT}}_{1}$ $32\times 32$ ImageNet ViT $16$ $192$ $1024$ ${\mathcal{D}}{\mathcal{T}}^{\text{LM}}_{1}$ LM1B Transformer LM $16$ $192$ $128$ $64$
F.2 Extended results
((a))
((b))
((c))
((d))
Figure 9: Evaluating generalization to wider networks. We train MLPs of different sizes to classify $32\times 32\times 3$ ImageNet images. Our optimizers are meta-trained for $1000$ inner steps (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We observe that $\mu\text{LO}_{M}$ generalize smoothly to all tasks, even surpassing the performance of VeLO (Fig.(c),(d)). We plot average training loss over $5$ seeds with standard error bars.
((a))
((b))
((c))
((d))
Figure 10: Meta-generalization to Cifar-10 and wider networks. We train MLPs of different sizes to classify Cifar-10 images. Our optimizers are meta-trained for $1000$ inner steps (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We observe that $\mu\text{LO}_{M}$ generalizes seamlessly to all tasks, reaching 00 loss before VeLO in subplots (b),(c),(d) while all other optimizers struggle to train stably. We plot average training loss over $5$ seeds with standard error bars.
Appendix G Coordinate evolution of MLP layers in $\mu$P for different optimizers

### v2

@@ 4.1.2 Meta-generalization to wider networks
Figure 5: Evaluating generalization capabilities of $\mu$LOs to deeper networks. The figures report the performance of learned optimizers for training depth-16 ViTs for image classification, Transformers for language modeling, and MLPs for image classification. We plot average training loss over $5$ seeds with standard error bars. In each case, $\mu$LOs show improved generalization and performance when compared to their SP counterparts.
4.1.3 Meta-generalization to deeper networks
In this section, we evaluate LO meta-generalization to deeper networks. Specifically, we increase the number of layers used in MLP, ViT, and LM tasks from $3$ to $16$, while being sure to select models that have widths within the meta-training range ($128-1024$) to avoid confounding the results. Figure 5 reports the performance of our multi-task learned optimizers on deeper networks. We observe that both $\mu\text{LO}_{M}$ and $\mu\text{VeLO}_{M}$ optimize stably throughout and generally outperform their counterparts, $\text{LO}_{M}$ and $\text{VeLO}_{M}$, by the end of training on each task, despite being meta-trained on MLPs of exactly the same depth. Moreover, $\text{LO}_{M}$ immediately diverges when optimizing the deep MLP while $\mu\text{LO}_{M}$ experience no instability. Similarly, $\text{VeLO}_{M}$ diverges on ViTs and Transformers, while $\mu\text{VeLO}_{M}$ performs well, especially on ViTs. This is remarkable as, unlike width, there is no theoretical justification for $\mu$P’s benefit to deeper networks. We hypothesize that $\mu$P’s stabilizing effect on the optimizee’s activations leads to this improvement generalization.
In summary, we find, empirically, that using $\mu$P during meta-training benefits the generalization of learned optimizers, including VeLO, to deeper networks.
@@ 4.1.3 Meta-generalization to deeper networks
Figure 6: Evaluating generalization capabilities of $\mu$LOs to longer training horizons. All optimizers are meta-trained for $1000$ steps of training (dotted red line), therefore, any optimization beyond $1000$ steps is considered out-of-distribution. We observe that $\mu$LOs seamlessly generalize to training horizons 25$\times$ longer than meta-training. In contrast, the best performing SP LO fails to decrease training loss (a), decreases it but suffers instabilities (b), or diverges after $8000$ steps (c).
@@ 4.1.4 Meta-generalization to longer training horizons
In this subsection, we evaluate the capability of $\mu$LOs to generalize to much longer training horizons than those seen during meta-training. Specifically, we use $\mu\text{LO}_{M}$ and $\text{LO}_{M}$ as well as $\mu\text{VeLO}_{M}$ and $\text{VeLO}_{M}$ to train three networks with width $w=1024$: a 3-layer MLP, ViT on $32\times 32\times 3$ ImageNet and a 3-layer Transformer for autoregressive language modeling on LM1B. Each model is trained for $25,000$ steps (25$\times$ the longest unroll seen at meta-training time). Figure 6 reports the training loss averaged over $5$ random seeds. We observe that $\mu\text{LO}_{M}$ and $\mu\text{VeLO}_{M}$ stably decrease training loss over time for each task, while $\text{LO}_{M}$ and $\text{VeLO}_{M}$ fail to decrease training loss (a), decreases it but suffers instabilities (b), or diverges after $8000$ steps (c). These results suggest that generalization to longer training horizons is another benefit of using $\mu$P with learned optimizers.
@@ 5 Limitations
While we have conducted a systematic empirical study and shown strong results within the scope of our study, there are some of limitations of our work. Specifically, (1) we do not meta-train on tasks other than MLPs for image classification and we do not provide evaluation of models wider than $8192$ for MLPs and $3072$/$12288$ (hidden/FFN size) for transformers due to computational constraints in our academic environment.
@@ 6 Conclusion
We have demonstrated that applying or $\mu$LO meta-training recipe produces optimizers with substantially improved meta-generalization properties when compared to strong baselines from previous work. Remarkably, our $\mu$LOs even surpass VeLO-4000 (meta-trained for $4000$ TPU months) on wider versions of in-distribution tasks. Moreover, our experiments also show that $\mu$LOs meta-trained with our recipe generalize better to wider and deeper out-of-distribution tasks than their SP counterparts. Moreover, when evaluated on much longer training tasks, we observe that $\mu$LOs have a stabilizing effect, enabling meta-generalization to much longer unrolls (25×25$\times$25 × maximum meta-training unroll length). All of the aforementioned benefits of $\mu$LOs come at zero extra computational cost compared to SP LOs. Our results outline a promising path forward for low-cost meta-training of learned optimizers that can generalize to large unseen tasks.
@@ Acknowledgements
Appendix A Proof of proposition 1
@@
The update produced by $f_{\phi}$ is denoted $\Delta W$ and we write $\nabla W$ the corresponding gradient, so that $\Delta W=f_{\phi}(\nabla W)$. For the sake of simplicity, $n$ will be the output size and $d$ the feature input size of our neural network. Our goal is to satisfy the desiderata of [Yang et al., 2022, Appendix J.2]. We assume our initialization follows Initialization-$\mu$ in Sec 3. Overall, our goal is to study strategy so that if $x_{i}=\Theta(1)$, then one needs to renormalize/initialize so that $(Wx)_{i}=\Theta(1)$ while $((W+\Delta W)x)_{i}=\Theta(1)$ so that the update is as large as possible. Note that given the assumptions on $f$, if $x=\Theta(\frac{1}{n})$, then $f(x)=\Theta(1)$.
@@
Appendix B Hand Designed Optimizer Hyperparameter Tuning
B.1 Tuning $\mu$Adam
We tune the $\mu$Adam baseline on the largest meta-training seen by our learned optimizers.$\mu$AdamM was, therefore, tuned using a $1024$ width MLP for $32\times 32\times 3$ ImageNet classification. As mentioned in section 4, we tune the learning rate and three multipliers: input multiplier, output multiplier, and the hidden learning rate multiplier. These multipliers correspond to adding a tunable constant to the pre-activation multiplier for input weights, the pre-activation multiplier for output weights, and the Adam LR for hidden weights (e.g., in Table 8 of Yang et al. [2022]). Specifically, we search for the learning rate in $\{0.1,0.01,0.001,0.0001\}$ and for each multiplier in $\{2^{-4},2^{-2},1,2^{2},2^{4}\}$. This results in a grid search of $500$ configurations, whose optimal values are reported in table 3.
Table 3: Best hyperparameters values for $\mu$Adam baseline. $\mu$Adam is tuned to optimize 3-layer W$=1024$ MLP for $32\times 32\times 3$ ImageNet classification.
@@ B.1 Tuning $\mu$Adam
B.2 Tuning AdamW
We tune the AdamW baseline on the largest meta-training seen by our learned optimizers. AdamW was, therefore, tuned using a $1024$ width MLP for $32\times 32\times 3$ ImageNet classification. As mentioned in section 4, we tune the learning rate, betas, and weight decay: LR, $\beta_{1}$, $\beta_{2}$, and the weights decay. Specifically, we search over the values of each hyperparameter reported in Table 4. This results in a grid search of $500$ configurations, whose optimal values are reported in table 5.
Table 4: Optimal Hyperparameters Found AdamW. Similar to the $\mu$Adam baseline, we tune all optimizers on a 3-layer W$=1024$ MLP ImageNet classification task and use a budget of approximately $500$ total runs.
Optimizer LR $\beta_{1}$ $\beta_{2}$ weight decay Total runs SP AdamWM Log Sample 14 from $[10^{-5},0.1]$$0.9$,$0.95$,$0.99$} {$0.95$,$0.99$,$0.999$} {$0.1$,$0.01$,$0.001$,$0.0001$} 504
Table 5: Optimal Hyperparameters Found AdamW. Similar to the $\mu$Adam baseline, we tune all optimizers on a 3-layer W$=1024$ MLP ImageNet classification task and use a budget of approximately $500$ total runs.
Optimizer LR $\beta_{1}$ $\beta_{2}$ weight decay Total runs SP AdamWM 0.000702 0.9 0.95 0.0001 504
Appendix C Meta-training with $\mu$LOs
General meta-training setup for small_fc_lopt
@@ Appendix C Meta-training with $\mu$LOs
Each small_fc_lopt [Metz et al., 2022a] learned optimizer is meta-trained for $5000$ steps of gradient descent using AdamW [Loshchilov and Hutter, 2019] and a linear warmup and cosine annealing schedule. We using PES [Vicol et al., 2021] to estimate meta-gradients with a truncation length of $50$ steps and sampling $8$ perturbations per task at each step with standard deviation $0.01$. For the inner optimization task, we used a maximum unroll length of $1000$ iterations; that is, all our learned optimizers see at most $1000$ steps of the inner optimization problem during meta-training. Unlike with $\mu$Adam, we do not tune the $\mu$P multipliers when meta-training $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$, instead, we set the all to $1$. All optimizers are meta-trained on a single A6000 GPU. $\mu\text{LO}_{S}$ and $\text{LO}_{S}$ take $8$ hours each to meta-train, while $\mu\text{LO}_{M}$ and $\text{LO}_{M}$ take $103$ hours.
General meta-training setup for VeLO
Each VeLO [Metz et al., 2022a] learned optimizer is meta-trained for $45000$ steps of gradient descent using AdamW [Loshchilov and Hutter, 2019] and a linear warmup and cosine annealing schedule. We using PES [Vicol et al., 2021] to estimate meta-gradients with a truncation length of $20$ steps and sampling $8$ perturbations per task at each step with standard deviation $0.01$. For the inner optimization task, we used a maximum unroll length of $1000$ iterations; that is, all our learned optimizers see at most $1000$ steps of the inner optimization problem during meta-training. Unlike with $\mu$Adam, we do not tune the $\mu$P multipliers when meta-training $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$, instead, we set the all to $1$. All optimizers are meta-trained on a single A6000 GPU. $\mu$VeLOMand VeLOM take $250$ hours to meta-train.
Meta-training hyperparameters for small_fc_lopt in $\mu$P
@@
While there are very few differences between $\mu$LOs and SP LOs, the effective step size for hidden layers is changed (see eq. 3) which could alter the optimal meta-training hyperparameters. Consequently, we conduct an ablation study on hyper-parameters choices for $\mu\text{LO}_{S}$. Specifically, using AdamW and gradient clipping with a linear warmup and cosine annealing LR schedule, we meta-train $\mu\text{LO}_{S}$ to train 3-layer width 128 MLPs to classify $32\times 32\times 3$ ImageNet Images. By default, we warmup linearly for 100 steps to a maximum learning rate of $33e-$ 3 and anneal the learning rate for $4900$ steps to a value of $31e-$ 3 with $\lambda_{1}=0.001$ (from equation 3) and sampling $8$ perturbations per step in PESVicol et al. [2021]. The above ablation varies the maximum learning rate $\in\{1e-2,3e-3,1e-3\}$ (always using $100$ steps of warmup and decaying to 0.3×0.3$\times$0.3 ×MaxLR), $\lambda_{1}\in\{0.001,0.01,0.1\}$, the number of steps (5k or 10k), and the number of perturbations (8 or 16). We observe that using all default values except for $\lambda_{1}=0.01$ yields one of the best solutions while being fast to train and stable during meta-training. We, therefore, select these hyperparameters to meta-train $\mu\text{LO}_{S}$ and $\mu\text{LO}_{M}$.
Meta-training hyperparameters for VeLO in $\mu$P
Unlike small_fc_lopt, we do not find it necessary to $\lambda_{1}$ from its default value. However, we do remove the multiplication by the current parameter norm used in the update equation to VeLO as it causes meta-training problems when initializing tensors to zero.
@@
While we use the same $\mu$P at meta-training and testing time, it is important to consider meta-training tasks that have similar training trajectories to their infinite width counterparts. In [Yang et al., 2022], authors provide discussions of these points for zero-shot hyperparameter transfer. Two notable guidelines are to initialize the output weight matrix to zero (as it will approach zero in the limit) and to use a relatively large key size when meta-training transformers. For all our tasks, we initialize the network’s final layer to zeros. While we do not meta-train on transformers, we suspect that the aforementioned transformer-specific guidelines may be useful.
Figure 7: Ablating Meta-training Hyperparameter for $\mu\text{LO}_{S}$. Using AdamW with a linear warmup and cosine annealing schedule, we meta-train $\mu\text{LO}_{S}$ to train 3-layer width 128 MLPs for classifying $32\times 32\times 3$ ImageNet Images. By default, we warmup linearly for 100 steps to a maximum learning rate of $33e-$ 3 and anneal the learning rate for $4900$ steps to a value of $31e-$ 3 with $\lambda_{1}=0.001$ (from equation 3) and sampling $8$ perturbations per step in PESVicol et al. [2021]. The above ablation varies the maximum learning rate $\in\{1e-2,3e-3,1e-3\}$ (always using $100$ steps of warmup and decaying to 0.3×0.3$\times$0.3 ×MaxLR), $\lambda_{1}\in\{0.001,0.01,0.1\}$, the number of steps (5k or 10k), and the number of perturbations (8 or 16). We observe that using all default values except for $\lambda_{1}=0.01$ yields one of the best solutions while being fast to train and stable during meta-training.
Appendix D Features of the learned optimizer
Table 6: Per-parameter features used as input to our learned optimizers. All the coefficients, $\beta_{i}$, are learnable parameters adjusted during meta-optimization. We replicate the table of [Joseph et al., 2023] for convenience.
@@ 3 $m$ Adafactor normalized values $m_{t,i}\times\textsc{row factor}\times\textsc{column factor}$
Appendix E List of Meta-testing Tasks
Table 7 reports the configuration of different testing tasks used to evaluate our optimizers. We note that we do not augment the ImageNet datasets we use in any way except for normalizing the images. We tokenize LM1B using a sentence piece tokenizer[Kudo and Richardson, 2018] with 32k vocabulary size. All evaluation tasks are run on A6000 $48$BG or A100 $80$GB GPUs for $5$ random seeds.
@@ Appendix E List of Meta-testing Tasks
[…]
Appendix F Coordinate evolution of MLP layers in $\mu$P for different optimizers
The following section presents the continuation of our experiments comparing pre-activation growth during training for SP LOs and $\mu$LOs with different meta-trainnig recipes, SP adam, and $\mu$Adam.